PEKKA K. SINERVO, Department of Physics, University of Toronto. The Status of the Top Quark Search at the Fermilab Tevatron^{1} The CDF and D0 Collaborations have completed analysis of approximately 20 pb^{−1} of proton-antiproton collisions at a centre-of-mass energy of 1.8 TeV. The data has been sifted for evidence of top quark production. The results of these searches will be presented, and prospects for future top quark physics at the Tevatron will be discussed.

1Supported by the Natural Sciences and Engineering Research Council of Canada